# Diabatic error and propagation of Majorana zero modes in interacting quantum dots systems

### v1

@@ III.2 Effect of nearest-neighbor Coulomb interactions
In this subsection, we explore the effect of repulsive nearest-neighbor Coulomb interactions, $V$, on the movement of the right-edge MZM, $\gamma_{2}$, in a moderately sized quantum-dot array with $L=16$. For very small quantum dot systems, the nearest-neighbor Coulomb interactions are highly detrimental to the stability of MZMs, as the finite $V$ leads to an overlap between the MZMs Bpandey2 . In realistic scenarios, the Coulomb interactions between neighboring quantum dots are always present. However, its influence in moderately sized quantum dot chains is not significant. This is because the nearest-neighbor Coulomb interactions, $V$, can be screened and minimized by the central hybrid superconducting region situated between neighboring quantum dots Tsintzis .
@@
Figure 3: (a) Site-dependent local density-of-states $LDOS(\omega,j)$ vs $j$ at $t=0$ and for different values of the Coulomb interaction $V$ using a system size $L=16$. For smaller values of $V\leq 0.1$, the MZMs are localized sharply on the edges. With an increase in $V$, the MZMs are localized exponentially, namely spread over a few more sites. The inset shows the $LDOS(\omega)$ vs $\omega$ for site $j=16$ and different values of $V$. The diabatic error $f(t)$ vs $t/\tau$ for different values of $\tau$ and for Coulomb interactions are shown in (b) $V$=0.1, (c) $V$=0.5, and (d) $V$=1.0. The diabatic error increases with increase in $V$ and for smaller values of $\tau$, namely for fast movement, as intuition indicates.
@@ III.3 Effect of disorder in hopping and pairing terms
where $\sigma$ is the standard deviation and $a$ is the mean of the distribution. We assume that the onsite chemical potential at time $t=0$ is zero for each site, achievable in quantum dot experiments by tuning the gate potentials attached to individual quantum dots ChunX . For the calculations, we choose the distribution to be centered around zero ($a=0$) with different standard deviations $\sigma$ up to $0.6$. For the static case, the results in Ref. Jay demonstrate that even a small number of quantum dots is sufficient to support a robust pair of localized MZMs under moderate disorder, in contrast to nanowire systems. We then study the effect of disorder on the movement of the right-edge MZM, $\gamma_{2}$, for a moderately sized quantum dot array with $L=16$.
@@
Figure 4(b) shows the disorder-average diabatic error $f(t)$ acquired by the system during the movement of the MZM $\gamma_{2}$ from site $j=16$ to site $j=10$ for different values of $\sigma$ up to $0.6$ and a fixed $\tau=36$. We find that the diabatic error slowly increases with time for all values of $\sigma$, but for weak disorder, such as $\sigma=0.2$, $f(t)$ is very close to the pristine system ($\sigma=0.0)$. However, for stronger disorder, such as $\sigma=0.6$, we found a sudden increase in the diabatic error. This is due to the fact that a strong deviation of $t_{h}^{k}/\Delta_{k}$ from $1$ reduces the many-body gap and makes it easier to have low-energy excitations as we try to move the MZM. In Fig. 4(c) we show the $LDOS(\omega,j,t)$ for $j=10$ at $t/\tau=6$ and $\sigma=0.2$ for a particular realization. We found sharp electron and hole peaks of equal height, confirming the movement of the MZM from $j=16$ to $j=10$ in a time $t/\tau=6$.
@@
Figure 4: (a) Site-dependent disorder averaged local density-of-states $LDOS(\omega,j)$ vs $j$ at time $t=0$ for different values of disorder strength $\sigma$ introduced in the hoppings $t^{h}_{k}$ and in the pairing terms $\Delta_{k}$. Inset shows the $LDOS(\omega,j)$ vs $\omega$ at site $j=16$ and for $\sigma=0.6$. The equal-height peaks of the electron and hole parts of the $LDOS(\omega)$ illustrate the existence of MZMs in the system for $\sigma\leq 0.6$ at $t=0$. (b) The disorder averaged diabatic error $f(t)$ in the system during movement of the right MZM $\gamma_{2}$ from site $j=16\to 10$, for different values of $\sigma$. Clearly, by increasing $\sigma$ the diabatic error slowly increases with time $t/\tau$. For $\sigma=0.6$ the diabatic error takes a larger jump and the system displays low-quality MZMs. (c) Electron and hole parts of the time-dependent $LDOS(\omega,j,t)$ at site $j=10$ for $\sigma=0.2$ for a particular value of the disorder realization (taken from a Gaussian distribution).
@@ III.4 Movement of the MZMs by Tuning Multiple Gates Simultaneously
Figure 5: Schematic of the movement of the right-edge MZM $\gamma_{2}$ by simultaneously tuning three gates in a six-quantum-dot system. (a) At time $t=0$, the system contains two MZMs ($\gamma_{1}$, $\gamma_{2}$), located at the left and right edges of the system, when working at the sweet-spot $t_{h}=\Delta$. (b) At time $t=T$, the right-edge MZM has moved to site $j=3$ using a simultaneous tuning of the time-dependent onsite chemical potentials, $\mu(t)$, applied to individual quantum dots. (c) Time-dependent local density-of-states, $LDOS(\omega,j,t)$, for different values of $t/\tau$ using $\tau=36$. (d) Diabatic error, $f(t)$, for the simultaneous movement of three gate potentials as a function of $t/\tau$ and for different values of $\tau$. (e) Diabatic error, $f(t)$, for the sequential movement of the three gate potentials as a function of $t/\tau$ and for different values of $\tau$.
In quantum-dot experiments, the gate attached to an individual quantum dot can be tuned either sequentially or simultaneously across multiple gates Dvir ; ChunX . In this subsection, we compare the diabatic error during the movement of the right-edge MZM, $\gamma_{2}$, for two scenarios: (i) when multiple gates are tuned simultaneously, and (ii) when the gates are tuned sequentially, one by one. To illustrate this, we consider a setup involving six quantum dots to perform the movement of the right-edge MZM, $\gamma_{2}$ (see Fig. 5(a)), which is experimentally feasible in current quantum-dot setups. We then tune the chemical potential, $\mu(t)=\mu_{0}t/\tau$ (where $\mu_{0}=12$), of the three rightmost quantum dots simultaneously (see Fig. 5(b)) at a quench rate of $1/\tau$. We chose very small values of repulsive Coulomb interaction $V=0.001$, to prevent the overlap of MZMs in the small system size $L=6$ that we consider for our numerical simulation. In Fig. 5(c), we present the time-dependent, site-dependent local density-of-states, $LDOS(\omega,j,t)$, for different values of $t/\tau$, when the three gate potentials, $\mu(t)$, are tuned simultaneously with $\tau=36$. Interestingly, for smaller values of $t/\tau$, the local density-of-states exhibits delocalized peaks spanning three sites, with a higher weight at site $j=3$. At the final time, $t/\tau=1$, the $LDOS(\omega,j,t)$ displays two localized peaks on sites $j=1$ and $j=3$, indicating that the right-edge MZM, $\gamma_{2}$, has been successfully transferred from $j=6$ to $j=3$.
@@
The above results demonstrate that the movement of the MZM in quantum-dot systems for quantum information processing can be completed well within the qubit lifetime ChunX , ensuring that it remains robust against decoherence and poisoning during the process. We also conclude that a combination of simultaneous and sequential movement of MZMs will be optimal for quantum information processing (braiding or fusion) in a network of quantum dots for topological quantum computations. The simultaneous movement of MZMs leads to delocalization over a larger number of sites, whereas the sequential movement provides better control and well-localized behavior of MZMs.
@@ IV Conclusion and Discussion
In the presence of nearest-neighbor repulsive Coulomb interactions, minimizing diabatic errors requires slower quench rates due to the effective gap reduction and increase in coherence length with stronger interactions. Interestingly, we find that moderate levels of disorder in hopping and pairing terms ($\sigma\leq 4$) do not significantly impact the diabatic errors. However, with further increase in disorder, we observe a sharp increase in diabatic errors, signifying MZM quality degradation caused by increased coherence length and reduced effective gap. We also observed similar trends in diabatic error for the movement of MZMs, in case of onsite quench disorder.
Our comparative study revealed that simultaneous tuning of gate potentials outperforms sequential tuning for slower quench rates, making it advantageous for efficient MZM movement. However, in quantum-dot networks designed for braiding experiments, a combination of simultaneous and sequential tuning offers greater control and better preserves the localized nature of MZMs. Such controlled movements are essential for minimizing overlaps with other MZMs, ensuring optimized quantum information processing in quantum-dot networks using MZMs.
Finally, using just six quantum-dots, we estimated the time scale to move the MZM form end sites to third sites. We show that our predicted time scale (0.3ns to 0.9ns) to move MZM in realistic setups is well within the qubit lifetime (1ms). We strongly believe that our novel theoretical results of diabatic error for the movement of Majoranas in the one-dimensional chain will be helpful for fusion and braiding experiments in the recently developed quantum-dot setups. Our predictions, based on time-dependent LDOS and fidelity measurements, align with current experimental capabilities, paving the way for future advancements in quantum-dot-based topological quantum computation.
@@ Acknowledgements.
The work of BP, SO, and ED was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division. The work of GA was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. The work of GKG was supported by the Texas Center for Superconductivity, University of Houston.
@@ Appendix A Non-analytic behavior of many-body gap
Figure 6: Difference in many-body energy levels $E_{3}$ (first excited stated above the gap) and $E_{2}$ (degenerate ground state) with change in sequential onsite chemical potential $\mu_{i}$ (from site $j=12\to 6$) as a parameter ($\mu_{seq}=\sum_{i}\mu_{i}$). For the numerical calculation, we used full diagonalization to find the many-body energy spectrum for $L=12$ sites, working at $V=0.05and\Delta=t_{h}=1$.
@@ Appendix B Disorder in on-site potential
Figure 7: Disorder averaged local density-of-states $LDOS(\omega,j)$ vs $j$ at time $t=0$ for different values of disorder strength $\sigma$, introduced in the onsite chemical potential $\mu_{j}$, and $t_{h}=\Delta=1$, $V=1.0$, $L=32$, $\mu_{0}=12$. Inset shows electron and hole parts of the local density-of-states $LDOS(\omega)$ vs $\omega$ at $t=0$, $j=1$ and for $\sigma=0$. (c) Disordered averaged electron and hole parts of the local density-of-states $LDOS(\omega)$ vs $\omega$ at $t=0$, $j=1$ and for $\sigma=1.2$ (b) The disorder averaged diabatic error $f(t)$ in the system while moving the right MZM $\gamma_{2}$ from site $j=32\to 24$, for different values of $\sigma$ and with increase in time $t/\tau$.
Figure 7 shows the effect of disorder similar to Fig. 4 except for using on-site disorder. We study the movement of the right-edge MZM, $\gamma_{2}$, for a moderately sized quantum dot array with $L=32$. In Fig. 7 (a), we show the site-dependent $LDOS(\omega,j,t)$ at $t=0$ for $t_{h}=\Delta=1$ and $V=1.0$. We found MZMs still localized at both edges but they develop a tail in the bulk, which increases with the strength of disorder. In the inset, we show the $LDOS(\omega,j,t)$ at site $j=1$ and time $t=0$. We found equal height electron and hole peaks, confirming the presence of MZMs. In Fig. 7(b), we show the accumulation of diabatic error with time as we move the right-edge MZM, $\gamma_{2}$, from site $j=32\to 24$. For small and moderate values of disorder strength, namely for $\sigma=0.4$ and $0.8$, the effect of disorder is small. For larger disorder strengths, namely for $\sigma=1.0$ and $1.2$, the diabatic error increases fast. In Fig. 7(c) we show the $LDOS(\omega,j,t)$ for $j=1$, $t=0$, and $\sigma=1.2$ for a particular realization. We still find equal height peaks for electron and hole, confirming the presence of a MZM, but the movement of these MZMs acquire large diabatic errors due to low-quality MZMs. We also see finite $LDOS$ away from $\omega=0$ because disorder introduces localized states close to the SC coherence peaks. These results demonstrate that the diabatic error remains small up to moderate values of disorder and interactions. In contrast, for nanowire systems, the effective gap is expected to nearly close under moderate disorder strength and for system sizes comparable to quantum-dot systems. Meanwhile, the effective gap in quantum-dot systems remains robust under moderate disorder and interactions, resulting in lower diabatic errors compared to nanowire systems Pan .

### v2

@@ III.2 Effect of nearest-neighbor Coulomb interactions
In this subsection, we explore the effect of repulsive nearest-neighbor Coulomb interactions, $V$, on the movement of the right-edge MZM, $\gamma_{2}$, in a moderately sized quantum-dot array with $L=16$. For very small quantum dot systems, the nearest-neighbor Coulomb interactions are highly detrimental to the stability of MZMs, as the finite $V$ leads to an overlap between the MZMs Bpandey2 . In realistic scenarios, the Coulomb interactions between neighboring quantum dots should be always present. However, its influence in moderately sized quantum dot chains is not significant. This is because the nearest-neighbor Coulomb interactions, $V$, can be screened and minimized by the central hybrid superconducting region situated between neighboring quantum dots Tsintzis .
@@
Figure 4: (a) Site-dependent local density-of-states $LDOS(\omega,j)$ vs $j$ at $t=0$ and for different values of the Coulomb interaction $V$ using a system size $L=16$. For smaller values of $V\leq 0.1$, the MZMs are localized sharply on the edges. With an increase in $V$, the MZMs are localized exponentially, namely spread over a few more sites. The inset shows the $LDOS(\omega)$ vs $\omega$ for site $j=16$ and different values of $V$. The diabatic error $f(t)$ vs $t/\tau$ for different values of $\tau$ and for Coulomb interactions as shown in (b) $V=0.1$, (c) $V=0.5$, and (d) $V=1.0$. The diabatic error increases with increase in $V$ and for smaller values of $\tau$, namely for fast movement, as intuition indicates.
@@ III.3 Effect of disorder in hopping and pairing terms
where $\sigma$ is the standard deviation and $a$ is the mean of the distribution. We assume that the onsite chemical potential at time $t=0$ is zero for each site, achievable in quantum dot experiments by tuning the gate potentials attached to individual quantum dots ChunX . For the calculations, we choose the distribution to be centered around zero ($a=0$) with different standard deviations $\sigma$ up to $0.6$.
For the static case, the results in Ref. Jay demonstrate that even a small number of quantum dots is sufficient to support a robust pair of localized MZMs under moderate disorder, in contrast to nanowire systems. We then study the effect of disorder on the movement of the right-edge MZM, $\gamma_{2}$, for a moderately sized quantum dot array with $L=16$.
@@
Figure 5(b) shows the disorder-average diabatic error $f(t)$ acquired by the system during the movement of the MZM $\gamma_{2}$ from site $j=16$ to site $j=10$ for different values of $\sigma$ up to $0.6$ and a fixed $\tau=36$. We find that the diabatic error slowly increases with time for all values of $\sigma$, but for weak disorder, such as $\sigma=0.2$, $f(t)$ is very close to the pristine system ($\sigma=0.0)$. However, for stronger disorder, such as $\sigma=0.6$, we found a sudden increase in the diabatic error. This is due to the fact that a strong deviation of $t_{h}^{k}/\Delta_{k}$ from $1$ reduces the many-body gap and makes it easier to have low-energy excitations as we try to move the MZM. Additionally, as disorder increases, the localization length of MZMs grows (Fig. 5(a)), leading to greater overlap with the left MZM ($\gamma_{1}$) during the movement of the right MZM ($\gamma_{2}$) from site $j=16$ to site $j=10$. This overlap contributes to a sudden increase in diabatic error for $\sigma\geq 0.4$ Thomale3 ; Rachel . In Fig. 5(c) we show the $LDOS(\omega,j,t)$ for $j=10$ at $t/\tau=6$ and $\sigma=0.2$ for a particular realization. We found sharp electron and hole peaks of equal height, confirming the movement of the MZM from $j=16$ to $j=10$ in a time $t/\tau=6$.
@@
Figure 5: (a) Site-dependent disorder averaged local density-of-states $LDOS(\omega,j)$ vs $j$ at time $t=0$ for different values of disorder strength $\sigma$ introduced in the hoppings $t^{h}_{k}$ and in the pairing terms $\Delta_{k}$. Inset shows the $LDOS(\omega,j)$ vs $\omega$ at site $j=16$ and for $\sigma=0.6$. The equal-height peaks of the electron and hole parts of the $LDOS(\omega)$ illustrate the existence of MZMs in the system for $\sigma\leq 0.6$ at $t=0$. (b) The disorder averaged diabatic error $f(t)$ in the system during movement of the right MZM $\gamma_{2}$ from site $j=16\to 10$, for different values of $\sigma$. Clearly, by increasing $\sigma$ the diabatic error slowly increases with time $t/\tau$. For $\sigma=0.6$ the diabatic error takes a larger jump and the system may have low-quality MZMs. (c) Electron and hole parts of the time-dependent $LDOS(\omega,j,t)$ at site $j=10$ for $\sigma=0.2$ for a particular value of the disorder realization (taken from a Gaussian distribution).
@@ III.4 Movement of the MZMs by Tuning Multiple Gates Simultaneously
Figure 6: Schematic of the movement of the right-edge MZM, $\gamma_{2}$, by simultaneously tuning three gates in a six-quantum-dot system. (a) At time $t=0$, the system contains two MZMs ($\gamma_{1}$, $\gamma_{2}$), located at the left and right edges, when operating at the sweet spot $t_{h}=\Delta$. (b) At time $t=T$, the right-edge MZM has moved to site $j=3$ via simultaneous tuning of the time-dependent onsite chemical potentials, $\mu(t)$, applied to individual quantum dots. (c) Time-dependent local density of states, $LDOS(\omega,j,t)$, for different values of $t/\tau$ using $\tau=36$. (d,f) Diabatic error, $f(t)$, for the simultaneous movement of three gate potentials as a function of $t/\tau$ and for different values of $\tau$ using: (d) linear-ramp protocol, (f) sine-squared ramp protocol. (e,g) Diabatic error, $f(t)$, for the sequential movement of the three gate potentials as a function of $t/\tau$ and for different values of $\tau$ using: (e) linear-ramp protocol, (g) sine-squared ramp protocol.
In quantum-dot experiments, the gate attached to an individual quantum dot can be tuned either sequentially or simultaneously across multiple gates Dvir ; ChunX . In this subsection, we compare the diabatic error during the movement of the right-edge MZM, $\gamma_{2}$, for two scenarios: (i) when multiple gates are tuned simultaneously, and (ii) when the gates are tuned sequentially, one by one. To illustrate this, we consider a setup involving six quantum dots to perform the movement of the right-edge MZM, $\gamma_{2}$ (see Fig. 6(a)), which is experimentally feasible in current quantum-dot setups. We then tune the chemical potential, $\mu(t)=\mu_{0}t/\tau$ (where $\mu_{0}=12$), of the three rightmost quantum dots simultaneously (see Fig. 6(b)) at a quench rate of $1/\tau$. We chose very small values of repulsive Coulomb interaction $V=0.001$, to prevent the overlap of MZMs in the small system size $L=6$ that we consider for our numerical simulation. In Fig. 6(c), we present the time-dependent, site-dependent local density-of-states, $LDOS(\omega,j,t)$, for different values of $t/\tau$, when the three gate potentials, $\mu(t)$, are tuned simultaneously with $\tau=36$. Interestingly, for smaller values of $t/\tau$, the local density-of-states exhibits delocalized peaks spanning three sites, with a higher weight on site $j=3$. At the final time, $t/\tau=1$, the $LDOS(\omega,j,t)$ displays two localized peaks on sites $j=1$ and $j=3$, indicating that the right-edge MZM, $\gamma_{2}$, has been successfully transferred from site $j=6$ to $j=3$.
@@
In Figs. 6(f,g), we analyze the diabatic error using a smooth time-dependent onsite potential Truong ; Sekania , $\mu(t)=\mu_{0}\sin^{2}\left(\frac{\pi t}{2\tau}\right)$, for two cases: (i) when multiple gates are tuned simultaneously (Fig. 6(f)) and (ii) when the gates are tuned sequentially, one by one (Fig. 6(g)). For faster movement ($\tau=12,24$) with simultaneous tuning of the smooth time-dependent gate potential, the diabatic error is comparable to that of linear gate potential tuning. This is expected, as a large amount of energy is quenched in a short time. Interestingly, for sequential and smooth time-dependent gate tuning (Fig. 6(g)), the diabatic error remains significantly lower in the range $24\leq\tau\leq 48$ compared to all other cases of diabatic movement (Figs. 6(d,e,f)). This reduction occurs because sequential smooth tuning mitigates nonanalytic behavior at each gate-tuning event in comparison to linear gate potential tuning Garrido . Remarkably, for larger $\tau\geq 60$ (slower tuning), the simultaneous tuning of the smooth time-dependent gate potential (Fig. 6(f)) results in the lowest diabatic error ($\sim 10^{-6}$) over time.
@@
The above results demonstrate that the movement of MZMs in quantum-dot systems for quantum information processing can be completed well within known qubit lifetimes ChunX , ensuring robustness against decoherence and quasiparticle poisoning during the process. Furthermore, we conclude that an optimal strategy for quantum information processing (such as braiding or fusion) in a network of quantum dots involves a combination of simultaneous and sequential MZM movement using a smooth time-dependent gate potential. The simultaneous movement of MZMs enhances delocalization over multiple sites, whereas the sequential movement provides better control and preserves the localized nature of MZMs.
@@ IV Conclusion and Discussion
In the presence of nearest-neighbor repulsive Coulomb interactions, minimizing diabatic errors requires slower quench rates due to the effective gap reduction and increase in coherence length with stronger interactions. Interestingly, we find that moderate levels of disorder in hopping and pairing terms ($\sigma\leq 0.4$) do not significantly impact the diabatic errors. However, with further increase in disorder, we observe a sharp increase in diabatic errors, signifying MZM quality degradation caused by increased coherence length and reduced effective gap. We also observed a similar trends in diabatic error for the movment of MZMs, in case of onsite quench disorder.
Our comparative study using linear and smooth gate tuning functions revealed that smooth gate tuning generates less diabatic error compared to linear gate potential tuning during MZM movement. We also find that simultaneous smooth gate potential tuning outperforms sequential tuning for slower quench rates, making it advantageous for efficient MZM transport. However, in quantum dot networks designed for braiding experiments, a combination of simultaneous and sequential tuning offers greater control and better preserves the localized nature of MZMs. Such controlled movements are essential for minimizing overlaps with other MZMs, ensuring optimized quantum information processing in quantum dot networks using MZMs.
Finally, using just six quantum-dots, we estimated the time scale to move the MZM from end sites to third sites. We show that our predicted time scale ( 0.3ns to 0.9ns) to move MZM in realistic setups is well within the qubit lifetime ( 1ms). We strongly believe that our novel theoretical results of diabatic error for the movement of Majoranas in the one-dimensional chain will be helpful for fusion and braiding experiments in the recently developed quantum-dot setups. Our predictions, based on time-dependent LDOS and fidelity measurements, align with current experimental capabilities, paving the way for future advancements in quantum-dot-based topological quantum computation.
@@ Acknowledgements.
The work of BP, SO, and ED was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division. The work of G.A. was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. The work of G. K. G. was supported by the Texas Center for Superconductivity, University of Houston.
Note added: All the data and simulation code are available from the corresponding author upon reasonable request.
@@ Appendix A Non-analytic behavior of many-body gap
Figure 7: Difference in many-body energy levels $E_{3}$ (first excited stated above the gap) and $E_{2}$ (degenerate ground state) with change in sequential onsite chemical potential $\mu_{i}$ (from site $j=12\to 6$) as a parameter ($\mu_{seq}=\sum_{i}\mu_{i}$). For the numerical calculation, we use full diagonalization to find the many-body energy spectrum for $L=12$ sites and $V=0.05,\Delta=t_{h}=1$
@@ Appendix B Disorder in on-site potential
Figure 8: Disorder averaged local density-of-states $LDOS(\omega,j)$ vs $j$ at $t=0$ for different values of disorder strength $\sigma$, introduced in the onsite chemical potential $\mu_{j}$, and $t_{h}=\Delta=1$, $V=1.0$, $L=32$, $\mu_{0}=12$. Inset shows electron and hole parts of the local density-of-states $LDOS(\omega)$ vs $\omega$ at $t=0$, $j=1$ and for $\sigma=0$. (c) Disordered averaged electron and hole part of the local density-of-states $LDOS(\omega)$ vs $\omega$ at $t=0$, $j=1$ and for $\sigma=1.2$ (b) The disorder averaged diabatic error $f(t)$ in the system while moving the right MZM $\gamma_{2}$ from site $j=32\to 24$, for different values of $\sigma$ and with increase in time $t/\tau$.
Figure 8 shows the effect of disorder similar to Fig. 4 except for on-site disorder. We study the movement of the right-edge MZM, $\gamma_{2}$, for a moderately sized quantum dot array with $L=32$. In Fig. 8 (a), we show the site-dependent $LDOS(\omega,j,t)$ at $t=0$ for $t_{h}=\Delta=1$ and $V=1.0$. We found MZMs still localized at both edges but they develop a tail in the bulk, which increases with the strength of disorder. In the inset, we show the $LDOS(\omega,j,t)$ at site $j=1$ and time $t=0$. We find equal-height electron and hole peaks, confirming the presence of MZMs. In Fig 8(b), we show the accumulation of diabatic error with time as we move the right-edge MZM, $\gamma_{2}$, from site $j=32\to 24$. For small and moderate values of disorder strength, namely for $\sigma=0.4$ and $0.8$, the effect of disorder is small. For larger disorder strengths, namely for $\sigma=1.0$ and $1.2$, the diabatic error increases fast. In Fig. 8(c) we show the $LDOS(\omega,j,t)$ for $j=1$, $t=0$, and $\sigma=1.2$ for a particular realization. We still find equal-height peaks for electron and hole, confirming the presence of a MZM, but the movement of these MZMs acquire large diabatic errors due to low-quality MZMs. We also see finite $LDOS$ away from $\omega=0$ because disorder introduces localized states close to the superconducting coherence peaks. These results demonstrate that the diabatic error remains small up to moderate values of disorder and interactions. In contrast, for nanowire systems, the effective gap is expected to nearly close under moderate disorder strength and for system sizes comparable to quantum-dot systems. Meanwhile, the effective gap in quantum-dot systems remains robust under moderate disorder and interactions, resulting in lower diabatic errors compared to nanowire systems Pan .